# Armijo Line-search Can Make (Stochastic) Gradient Descent Provably Faster

### v2

@@ C.2 Helper Lemmas
[…]
@@
See 3
@@
Using the SGD update: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla{f_{t}}(\theta_{t})$, for the comparator $u$ s.t. $f(u)\leq f(\theta_{t})$ for all $t\in[T]$,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $=\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,\langle\nabla{f_{t}}(\theta_{t}),\theta_{t}-u\rangle+{\eta_{t}}^{2}\,\left\|\nabla{f_{t}}(\theta_{t})\right\|_{2}^{2}$ (33)
@@
Taking expectation w.r.t. the randomness at iteration $t$
$\implies\mathbb{E}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\mathbb{E}\left[{\eta_{t}}\underbrace{\left(2-\frac{1}{c}\right)}_{\text{Positive}}\,\underbrace{[{f_{t}}(\theta_{t})-{f_{t}}^{*}]}_{\text{Positive}}\right]+2\,\mathbb{E}\left[{\eta_{t}}\underbrace{[{f_{t}}(u)-{f_{t}}^{*}]}_{\text{Positive}}\right]$ (34)
From the line-search, we know that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{C^{\prime}}{{f_{t}}(\theta_{t})}\right\}$ where $C^{\prime}:=\frac{1}{\lambda_{1}}$.
Using this result, let us first upper-bound $-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$. For this we will consider two cases.
Case (1): If ${f_{t}}(\theta_{t})<f(\theta_{t})$,
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max}\,\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\,,C^{\prime}\,\mathbb{E}\left[\frac{{f_{t}}(\theta_{t})-{f_{t}}^{*}}{{f_{t}}(\theta_{t})}\right]\right\}$ (Lower-bound on ${\eta_{t}}$)
$\leq-\min\left\{\eta_{\max}\,\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right],C^{\prime}\,\mathbb{E}\left[\frac{{f_{t}}(\theta_{t})-{f_{t}}^{*}}{f(\theta_{t})}\right]\right\}$ (Using the case (1) condition)
$=-\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}\,\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]$
$=-\underbrace{\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}}_{\text{Positive}}\,\bigg{[}\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}(u)\right]+\underbrace{\mathbb{E}\left[{f_{t}}(u)-{f_{t}}^{*}\right]}_{\text{Positive}}\bigg{]}$ (Add/Subtract ${f_{t}}(u)$)
$\implies-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]$
Case (2): If ${f_{t}}(\theta_{t})\geq f(\theta_{t})$,
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-{f_{t}}^{*}\right]\right]$ (Using the case (2) condition)
$=-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-{f_{t}}(u)\right]\right]-\underbrace{\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(u)-{f_{t}}^{*}\right]\right]}_{\text{Positive}}$ (Add/subtract ${f_{t}}(u)$)
$\implies-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]-\mathbb{E}\left[{\eta_{t}}\left[f(u)-{f_{t}}(u)\right]\right]$ (Add/subtract $f(u)$)
Let us consider two sub-cases: Case (i): If $f(u)-{f_{t}}(u)\geq 0$:
@@
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]$
Case (ii): If $f(u)-{f_{t}}(u)<0$:
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]+\mathbb{E}\underbrace{\left[{\eta_{t}}\left[{f_{t}}(u)-f(u)\right]\right]}_{\text{Positive}}$
$\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]+\eta_{\max}\,\mathbb{E}\underbrace{\left[{f_{t}}(u)-f(u)\right]}_{=0}$
$\implies-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]$
Hence, in both sub-cases, we get that,
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\underbrace{\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]}_{\text{Positive}}\leq-\min\left\{\eta_{\max}\left[f(\theta_{t})-f(u)\right],C^{\prime}\,\mathbb{E}\left[\frac{f(\theta_{t})-f(u)}{{f_{t}}(\theta_{t})}\right]\right\}$ (Using the lower-bound on ${\eta_{t}}$)
@@
$=-\min\left\{\eta_{\max}\left[f(\theta_{t})-f(u)\right],C^{\prime}\,[f(\theta_{t})-f(u)]\,\mathbb{E}\left[\frac{1}{{f_{t}}(\theta_{t})}\right]\right\}$ (Since $f(\theta_{t})-f(u)$ is independent of the randomness at iteration $t$)
$\leq-\min\left\{\eta_{\max}\left[f(\theta_{t})-f(u)\right],C^{\prime}\,[f(\theta_{t})-f(u)]\,\left[\frac{1}{\mathbb{E}[{f_{t}}(\theta_{t})]}\right]\right\}$ (Jensen’s inequality since $1/x$ is convex)
$\implies-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max}\left[f(\theta_{t})-f(u)\right],C^{\prime}\,\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]\right\}$ (Unbiasedness)
Hence, in both cases, we get that,
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]$
Combining the above relations,
$\mathbb{E}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]+2\mathbb{E}\left[{\eta_{t}}\underbrace{[{f_{t}}(u)-{f_{t}}^{*}]}_{>0}\right]$ (where $C:=C^{\prime}\,\frac{2c-1}{c}$)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]+2\eta_{\max}\,\underbrace{\mathbb{E}[{f_{t}}(u)-{f_{t}}^{*}]}_{:=\chi^{2}(u)}$
@@
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]+2\eta_{\max}\,\chi^{2}(u)$
Taking an expectation w.r.t the randomness iterations $t=0$ to $T-1$, and recursing, we conclude that
@@
$\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,\chi^{2}(u)\,T$
@@
For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{\lambda_{\max}[X^{T}X]}{4n}\leq 1$, satisfies Assn. 2 with $L_{0}=0$, $L_{1}=8$ and Assn. 3 with $\nu=8$, $\omega=0$.
@@
Consider $u=\beta u^{*}$ where $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon^{2}}\right)$ implies that $f(u)\leq\epsilon^{2}$. Since $\epsilon\leq 1$, $f(u)\leq\epsilon$.
For the $\epsilon$ and $T$ defined in the theorem statement, define the event $G=\{\inf_{t\in[T]}[f(\theta_{t})-f(u)]\leq\epsilon\}$. Using the law of total expectation,
$\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]\right]$ $=\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]|G\right]\,\Pr[G]+\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]|G^{c}\right]\,\Pr[G^{c}]$ (37)
$\leq\epsilon+\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]|G^{c}\right]$ (By definition of $G$)
Hence, we need to bound $\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]\right]$ conditioned on the event $G^{c}=\{\inf_{t\in[T]}[f(\theta_{t})-f(u)]\geq\epsilon\}$.
Conditioned on $G^{c}$, we know that for all $t\in[T]$, $f(\theta_{t})-f(u)\geq\epsilon$. Since $f(u)\leq f(\theta_{t})$ for all $t\in[T]$, we can use Lemma 3 with the comparator $u=\beta u^{*}$ and obtain the following inequality for SGD-SLS with $\theta_{0}=0$, where $C=\frac{2c-1}{c}\,\frac{1}{\lambda_{1}}$
$\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\left\|u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,\chi^{2}(u)\,T$
$\leq\left\|u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,f(u)\,T$ (Since $\chi^{2}(u)\leq f(u)$)
Setting $\eta_{\max}=\frac{C}{\epsilon}$ ensures that, conditioned on $G^{c}$, $\eta_{\max}\geq\frac{C}{f(\theta_{t})}$ for all $t\leq T$. Hence,
$\leq\left\|u\right\|_{2}^{2}-C\,\sum_{t=0}^{T-1}\mathbb{E}\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]+\frac{2C\,f(u)}{\epsilon}\,T$
Since $f(u)\leq\epsilon^{2}\leq\epsilon$ and conditioned on $G^{c}$, $\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]\geq\frac{\epsilon}{f(u)+\epsilon}$,
$\leq\left\|u\right\|_{2}^{2}-C\,\sum_{t=0}^{T-1}\mathbb{E}\left[\frac{\epsilon}{f(u)+\epsilon}\right]+\frac{2C\,f(u)}{\epsilon}\,T$
$\leq\beta^{2}-\frac{C\,T}{2}+2C\,\epsilon\,T$ (Since $u=\beta u^{*}$ and $f(u)\leq\epsilon^{2}\leq\epsilon$)
$\implies\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\beta^{2}-\frac{C\,T}{4}$ (For $\epsilon\leq\frac{1}{8}$)

### v3

@@ C.2 Helper Lemmas
$T$ $>\max\big{\{}1,2^{5}\,3^{3}\,5^{2}\,K^{2}\big{\}}\left(\frac{f^{*}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-f^{*}}{\epsilon}\right)$ (since $\lambda_{1}=3\frac{L_{1}(\nu+1)}{1-c}$ and $c=\frac{1}{2}$)
@@
Lemma 5.
For a finite sum $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}$, if $\min f_{i}(\theta)=0$ for all $i$ and an arbitrary comparator $u$, then iteration $t$ of SGD-SLS satisfies the following bound:
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,{\eta_{t}}\,[{f_{t}}(\theta_{t})]+2{\eta_{t}}[{f_{t}}(u)]\,.$
@@
Using the SGD update: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla{f_{t}}(\theta_{t})$, for a comparator $u$,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $=\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,\langle\nabla{f_{t}}(\theta_{t}),\theta_{t}-u\rangle+{\eta_{t}}^{2}\,\left\|\nabla{f_{t}}(\theta_{t})\right\|_{2}^{2}$
@@
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,{\eta_{t}}\,[{f_{t}}(\theta_{t})-{f_{t}}^{*}]+2{\eta_{t}}[{f_{t}}(u)-{f_{t}}^{*}]$
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,{\eta_{t}}\,[{f_{t}}(\theta_{t})]+2{\eta_{t}}[{f_{t}}(u)]$ (Since ${f_{t}}^{*}=0$)
See 4
Proof.
For the $\epsilon$ and $T$ defined in the theorem statement, define the event $G=\{\inf_{t\in[T]}[f(\theta_{t})]\leq\epsilon\}$. Using the law of total expectation,
$\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ $=\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G\right]\,\Pr[G]+\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G^{c}\right]\,\Pr[G^{c}]$ (28)
$\leq\epsilon+\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G^{c}\right]$ (By definition of $G$)
@@
Hence, we need to bound $\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ conditioned on the event $G^{c}=\{\inf_{t\in[T]}f(\theta_{t})\geq\epsilon\}$.
Conditioned on $G^{c}$: For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{\lambda_{\max}[X^{T}X]}{4n}\leq 1$. Using the standard guarantee for Armijo line-search, we know that $\eta_{\max}\geq\min\left\{\eta_{\max},\frac{2\,(1-c)}{L}\right\}$. Since $f_{t}(\theta)\geq 0$, combining this with Lemma 5,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\min\left\{\eta_{\max},\frac{2\,(1-c)}{L}\right\}{f_{t}}(\theta_{t})+2\eta_{\max}[{f_{t}}(u)]$
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{1}{2}\,\min\left\{\eta_{\max},\frac{2}{3}\right\}{f_{t}}(\theta_{t})+2\eta_{\max}[{f_{t}}(u)]$ (Since $L\leq 1$ and setting $c=\frac{2}{3}$)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{1}{3}\,{f_{t}}(\theta_{t})+\frac{2{f_{t}}(u)}{\epsilon}$ (Setting $\eta_{\max}=\frac{1}{\epsilon}$ and since $\epsilon\leq 1$)
Taking an expectation over the randomness in iteration $t$, conditioned on the past iterates,
@@
$\implies\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{f(\theta_{t})}{3}+\frac{2f(u)}{\epsilon}$ (Since $u$ is independent of the randomness at iteration $t$)
Consider $u=\beta u^{*}$ where $\beta=\frac{1}{\gamma}\,\ln\left(\frac{12}{\epsilon^{2}}\right)$ implies that $f(u)\leq\frac{\epsilon^{2}}{12}$. Since $\epsilon\leq 1$, $f(u)\leq\epsilon$. Using this relation with the above inequality, we get,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{f(\theta_{t})}{3}+\frac{\epsilon}{6}$
From the conditioning on $G^{c}$, we know that $f(\theta_{t})\geq\epsilon$. Hence,
$\implies\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\epsilon}{3}+\frac{\epsilon}{6}=\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\epsilon}{6}$
Taking expectation w.r.t. the randomness from iterations $t=0$ to T1T-1,
$\mathbb{E}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\mathbb{E}[\left\|\theta_{t}-u\right\|_{2}^{2}]-\frac{\epsilon}{6}$
@@
$\implies\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{\epsilon\,T}{6}$ (Summing from $t=0$ to T1T-1)
@@
$\leq\beta^{2}-\underbrace{\frac{\epsilon}{6}}_{:=C}\,T$ (Since $\theta_{0}=0$ and $\left\|u\right\|=\beta$)
$\mathbb{E}[f(\theta_{T})-f(u)]$ $\leq\frac{\epsilon}{2}+\underbrace{[\nu^{2}+1]}_{:=L}\,\mathbb{E}\frac{\left\|\theta_{T}-u\right\|_{2}^{2}}{2}\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left[\beta^{2}-C\,T\right]$
$\implies\mathbb{E}[f(\theta_{T})]$ $\leq\frac{3\epsilon}{2}+\frac{L}{2}\,\left[\beta^{2}-C\,T\right]$ (Since $f(u)\leq\epsilon$)
In order to ensure that $\mathbb{E}[f(\theta_{T})]\leq\frac{3\epsilon}{2}$, it is sufficient to set $T$ as:
$T\geq\frac{\beta^{2}}{C}$ $=\frac{6}{\epsilon\,\gamma^{2}}\,\left[\ln\left(\frac{12}{\epsilon^{2}}\right)\right]^{2}$
Since the above calculations were conditioned on $G^{c}$, we can conclude that after $T=\frac{6}{\epsilon\,\gamma^{2}}\,\left[\ln\left(\frac{12}{\epsilon^{2}}\right)\right]^{2}$ iterations, $\mathbb{E}[\inf_{t\in[T]}f(\theta_{T})|G^{c}]\leq\frac{3\epsilon}{2}$. Hence, from Eq. 29
$\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ $\leq\frac{5\epsilon}{2}$
@@
For the $\epsilon$ and $T$ defined in the theorem statement, define the event $G=\{\inf_{t\in[T]}[f(\theta_{t})]\leq\epsilon\}$. Using the law of total expectation,
$\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ $=\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G\right]\,\Pr[G]+\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G^{c}\right]\,\Pr[G^{c}]$ (29)
$\leq\epsilon+\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G^{c}\right]$ (By definition of $G$)
Hence, we need to bound $\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ conditioned on the event $G^{c}=\{\inf_{t\in[T]}f(\theta_{t})\geq\epsilon\}$.
Conditioned on $G^{c}$: Note that from Lemma 1, we know that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{C^{\prime}}{{f_{t}}(\theta_{t})}\right\}$ where $C^{\prime}:=\frac{1}{\lambda_{1}}$. Since $f_{t}(\theta)\geq 0$, combining this with Lemma 5, we get that,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}+2\eta_{\max}[{f_{t}}(u)]$
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\underbrace{\frac{\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}}_{:=(*)}+2\eta_{\max}[{f_{t}}(u)]$ (Setting $c=\frac{2}{3}$)
In order to simplify (*), we consider two cases that depend on whether or not ${f_{t}}(\theta_{t})\leq\epsilon$.
$(*)$ $=\frac{\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})\geq\epsilon\}+\frac{\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})<\epsilon\}$
$\implies(*)$ $\geq\frac{\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})\geq\epsilon\}\geq\frac{\min\left\{\eta_{\max}\epsilon,C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})\geq\epsilon\}$
Combining the above inequalities, we get that,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{\eta_{\max}\epsilon,C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})\geq\epsilon\}+2\eta_{\max}[{f_{t}}(u)]$
Taking an expectation over the randomness in iteration $t$, conditioned on the past iterates,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{\eta_{\max}\epsilon,C^{\prime}\right\}}{2}\,\Pr[f_{t}(\theta_{t})\geq\epsilon]+2\eta_{\max}[f(u)]$ (Since $\eta_{\max}$ and $u$ are both independent of the randomness at iteration $t$)
From the conditioning on $G^{c}$, we know that $f(\theta_{t})\geq\epsilon$. Since $f$ is the average of the $f_{i}$ losses, we know that there exists at least one $j$ s.t. $f_{j}(\theta_{t})\geq\epsilon$. Hence, conditioned on the past iterates, $\Pr[f_{t}(\theta_{t})]\geq\epsilon\geq\frac{1}{n}$. Combining this lower bound with the above inequality, we get that,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{\eta_{\max}\epsilon,C^{\prime}\right\}}{2n}+2\eta_{\max}[f(u)]$
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{1,C^{\prime}\right\}}{2n}+\frac{2\,f(u)}{\epsilon}$ (Setting $\eta_{\max}=\frac{1}{\epsilon}$)
@@
Consider $u=\beta u^{*}$ where $\beta=\frac{1}{\gamma}\,\ln\left(\frac{n}{\epsilon^{2}}\right)$ implies that $f(u)\leq\frac{\epsilon^{2}}{n}$. Since $\epsilon\leq 1$ and $n\geq 1$, $f(u)\leq\epsilon$. Using this relation with the above inequality, we get,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{1,C^{\prime}\right\}}{2n}+\frac{2\epsilon}{n}$
For $\epsilon\leq\frac{\min\left\{1,C^{\prime}\right\}}{8}\leq 1$,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{1,C^{\prime}\right\}}{4n}$
Taking expectation w.r.t. the randomness from iterations $t=0$ to T1T-1,
$\mathbb{E}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\mathbb{E}[\left\|\theta_{t}-u\right\|_{2}^{2}]-\frac{\min\left\{1,C^{\prime}\right\}}{4n}$
$\implies\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{\min\left\{1,C^{\prime}\right\}}{4n}\,T$ (Summing from $t=0$ to T1T-1)
$\leq\beta^{2}-\underbrace{\frac{\min\left\{1,C^{\prime}\right\}}{4n}}_{:=C}\,T$ (Since $\theta_{0}=0$ and $\left\|u\right\|=\beta$)